OpenStudy (anonymous):

MnO2 formula mass= molar mass= C5H12 formula mass= molar mass= XeF6 formula mass= molar mass=

OpenStudy (anonymous):

Mn = 54,93 O = 15,999 MnO2 = 54,93 + (2*15,999) = ??? g/mol C = 12,011 H = 1,008 C5H12=(5*12,011)+(12*1,008) = ??? g/mol Xe = 131,29 F = 18,998 XeF6 = 131,29 + (6* 18,998) = ??? g/mol

OpenStudy (anonymous):

so 86.91 formula mass and 86.936845 molar mass

OpenStudy (anonymous):

you wrote yourself formula mass= molar mass so it is for MnO2 = 86,928 g/mol

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

not a problem ;)